# SkinGrip: An Adaptive Soft Robotic Manipulator with Capacitive Sensing for Whole-Limb Bathing Assistance

### v1

@@ III-B End Effector Control from Capacitive Sensing
Baseline Rigid End Effector Control Similarly, for the baseline end effector, we employ a $2\times 3$ grid of electrodes, shown in Fig. 2 (g), to detect the relative pose of the end effector in relation to the human limb. This end effector follows the same design and control paradigm specified in [7, 9]. Specifically, the $2$-left-middle and $5$-right-middle electrodes are essential for identifying contact (along the $z$ axis) between the baseline end effector and human skin. The four electrodes, numbered $1,4,3$ and $6$, enable the detection of displacement along the $x$ axis and the measurement of the yaw rotation $\alpha$ on the surface of the limb, ensuring that the center line of the end effector remains aligned with the central axis of a human limb. Furthermore, the two rows of six electrodes collectively facilitate the measurement of roll rotation $\gamma$ on the limb’s surface, ensuring that the bottom surface of the end effector remains parallel to human skin. Note that we do not detect translations along the $y$ axis or rotation $\beta$ around the $y$-axis, as shown in Fig. 2 (f). The end effector translates at a fixed velocity of 4 cm/s along the axis of the limb to ensure bathing task progress.
@@
Algorithm 2 defines the capacitive servoing control strategy for the baseline rigid end effector. Similar to the approach used with the SkinGrip, at each time step, we normalize all real-time capacitance values to the range $[0,1]$ using the formula $\bm{C}_{norm}^{r}=\frac{\sum\bm{C}^{r}_{t-4:t}/n-\bm{C}_{min}^{r}}{\bm{C}_{max}^{r}-\bm{C}_{min}^{r}}$, where $\bm{C}_{max}^{r}$ and $\bm{C}_{min}^{r}\in\mathbb{R}^{6}$ denote the maximum and minimum capacitance values, respectively, for each electrode. We establish a threshold $\bm{C}_{th}^{r}\in\mathbb{R}^{6}$ to ensure “full contact” between the baseline end effector and the human skin. The control for the baseline end effector is then defined as:
@@
where $x$ and $z$ represent the x- and z-coordinate values of the displacements in the baseline end effector frame, and $\alpha$ and $\gamma$ denote the orientation values for roll and yaw displacements in the same frame (Fig. 2 (f)). The term $\bm{K}_{p}^{r}\in\mathbb{R}^{4\times 4}$ denotes a diagonal matrix for the proportional (P) gains. Specifically, $x_{r}=(\bm{C}_{norm}^{r})_{1}-(\bm{C}_{norm}^{r})_{3}+(\bm{C}_{norm}^{r})_{4}-(\bm{C}_{norm}^{r})_{6}$ is determined by the normalized capacitance values from electrodes in both the first row (electrodes $1$ and $3$) and the second row (electrodes $4$ and $6$), which are used to measure lateral displacement along the x-axis (Fig. 2 (g)). Similarly, $z_{r}=(\bm{C}_{th}^{r})_{2}-(\bm{C}_{norm}^{r})_{2}+(\bm{C}_{th}^{r})_{5}-(\bm{C}_{norm}^{r})_{5}$ is calculated from the difference between the threshold and normalized capacitance values of electrodes $2$ and $5$, enabling the assessment of vertical distance along the $z$-axis. The orientation angles $\alpha_{r}=(\bm{C}_{norm}^{r})_{4}-(\bm{C}_{norm}^{r})_{1}+(\bm{C}_{norm}^{r})_{3}-(\bm{C}_{norm}^{r})_{6}$ and $\gamma_{r}=(\bm{C}_{norm}^{r})_{1}+(\bm{C}_{norm}^{r})_{2}+(\bm{C}_{norm}^{r})_{3}-(\bm{C}_{norm}^{r})_{4}-(\bm{C}_{norm}^{r})_{5}-(\bm{C}_{norm}^{r})_{6}$ are obtained from the configuration of six electrodes in two rows, which facilitate the detection of roll and yaw movements relative to the limb’s surface and its central axis, respectively.
III-C SkinGrip Characterization
In order to understand how SkinGrip applies pressure when interacting with a human limb—an essential aspect of ensuring that the manipulator operates safely and effectively—we characterize the contact pressure applied across the surface of the soft fingers as the SkinGrip makes contact with and contracts around a human limb. To capture the ground truth contact pressure, we wrapped a participant’s limb in a fabric pressure-sensing mat (BodiTrak2 Pro [27]) before closing the manipulator around the wrapped limb. We repeat three trials of this procedure for the participant’s forearm and lower leg ($33$ cm and $38.5$ cm circumference, respectively), then crop the recorded pressure maps to each of the SkinGrip’s fingers. For each limb, we compute the average contact pressure map across trials and apply a Gaussian filter ($\sigma=1.4$) to smooth the resulting map for each finger. The setup for the experiments and the contact pressure mapping to each finger, when grasping the forearm and lower leg, are shown in Fig. 4. Note that due to limited sensitivity of the pressure sensing mat (minimum sensing threshold = 2.5kg) used in the experiments, minimum (zero) pressure does not imply that there is no contact.
When grasping either limb, the highest contact pressure is primarily observed at the tips of the manipulator’s fingers. Each finger has a second pressure cluster of lower amplitude, occurring at the center of the left fingers and the top (i.e. at the joint) of the right finger. Since the limbs are not truly cylindrical, the left finger wraps around a region of smaller diameter than the right finger, resulting in different contact pressure distributions across the finger on the left and right. These experimental results offer insights into how pressure is generally applied by the SkinGrip to the human body.
Figure 4: We measure the relative contact pressure applied by the SkinGrip to a pressure-sensing mat wrapped around the forearm and lower leg. We then map the recorded pressure to each of the soft fingers. Light and dark regions of the heatmap correspond to high and low pressure respectively. The darkest (zero) regions do not indicate that there is no contact.
@@ IV Study Design
To assess the cleaning performance of both our SkinGrip and the baseline end effector, we conducted a human-robot study involving $12$ able-bodied adult participants [$2$ females, $10$ males, age: $26.9$ ($6.5$) mean (SD) years, weight: $73.3$ ($9.3$) kg, height: $176.6$ ($8.2$) cm]. This study was approved by Carnegie Mellon University’s Institutional Review Board under 2022.00000273.
The experimental setup is illustrated in Fig. 5 (a). Participants lie on a hospital bed with the head and foot of the bed adjustable for comfort based on individual preferences. A limb holder secures the participant’s limb parallel to the ground plane during the cleaning task, and three cameras are positioned to capture top, side, and bottom views of the participant’s limbs. Prior to cleaning, three rings of shaving cream are applied to the limb that the robot must clean off to demonstrate skin coverage and bathing performance. Shaving cream was selected as it is skin-safe, straightforward to apply, and has a color that contrasts with skin tones to support visual analysis in cleaning performance. Recorded images from all three cameras are used to evaluate the cleaning performance of both the soft and baseline end effectors. An additional phone camera records video throughout the experiment. For each participant and each end effector, the robot conducted 2 bathing attempts for both the arm and the leg, resulting in 8 bathing trials per participant. We randomized ordering between the two end effectors for each participant.
At the beginning of an experiment with each participant, we established specific capacitance thresholds for both end effectors. The thresholds for the SkinGrip were identified at the point where participants reported comfort and confirmed “full contact” between the soft finger of the end effector and their limb as the soft fingers encircled their limb. In contrast, thresholds for the rigid end effector were set based on participant reports of comfortable pressure when the end effector, held at a neutral orientation with no roll, pitch, or yaw, was placed centrally against the skin of their limb.
@@ V-C Discussion
Through this research, we have illustrated the practicality and potential benefits of our soft robotic manipulator in enhancing the quality of bathing assistance, with an emphasis on safety, comfort, and effectiveness. However, our work builds on a few assumptions that could benefit from future advances. One key assumption is the expectation that an individual will be able to position their limb static, lifted off of a bed by a limb support. Second, we evaluated the proposed SkinGrip primarily on limb bathing scenarios. Although our SkinGrip has potential to clean other body parts, such as the back and torso, the control methods required for these tasks are still necessary. An additional future development may include the design of computer vision techniques to track regions of a limb that have not been fully cleaned in order to perform a targeted cleaning maneuver around those regions.
@@ VI Conclusion
In this work, we introduced an innovative soft deformable robotic manipulator—SkinGrip, designed for high adaptability and conforming to the contours of human limbs. By incorporating a capacitive servoing control scheme, the SkinGrip is able to detect contact and tightness of its soft fingers around the circumference of a human limb, while also effectively traversing the limb to perform robot-assisted bathing. We evaluate the SkinGrip and control strategy in a human experiment in which a mobile manipulator assists with cleaning off the arm and leg of 12 human participants. Our experiments highlight the SkinGrip and capacitive servoing control strategy’s ability to clean and adapt to various limb sizes, including differences in diameter and length. Notably, the SkinGrip achieves an overall average of $88.8\%$ cleaning effectiveness on the arm and $81.4\%$ on the leg, significantly outperforming a baseline rigid end effector design. Participant-provided feedback reinforces this finding, in which we observe a statistically significant difference in reported preferences for a soft manipulation strategy in terms of safety, comfort, and thorough skin bathing capabilities.

### v2

@@ III-B End Effector Operation from Capacitive Sensing
Next, we carried out tests to monitor the variations in capacitance values as the SkinGrip approaches a human limb and encloses the limb until full contact is established between the two soft fingers and the human skin. This process is illustrated in Fig. 3 (c) and (d). The end effector continues to approach the human limb with its fingertips open so long as the value of the $7$-left inner electrode (Fig. 2 (c)) falls below the threshold $C_{th}^{s2}$, indicating a lack of contact between the upper portion of the soft finger and the human skin. The control system adjusts the manipulator’s position $z_{t}$ in (2) to regulate the touch interaction with the skin.
When the value of the left inner sensor reaches the threshold $C_{th}^{s2}$, the two soft fingers commence a closure phase. During the closure phase the two fingers progressively close and surround the human limb for full contact. The closure phase ends when the sum of errors between the current and desired values of the six electrodes is minimized, indicating full contact. This is achieved by dynamically adjusting the motor to bend the fingers. The system then maintains this full contact dynamically, using a proportional controller (see 1) for continuous feedback and adjustment, and continuously adapting $z_{t}$ and $\gamma_{t}$ according with the P controller in (2). This ensures the fingers adapt to changes in limb position or size, maintaining contact by adjusting tendon tension based on feedback from all electrodes. Finally, the Stretch RE1 maneuvers the soft manipulator along the central axis of the human limb at a constant velocity of $4$ cm/s, starting from one end of the limb to the other end. In this process, the manipulator’s $z$ and $\gamma$ positions will be updated using 2, and the two soft fingers will be dynamically controlled using 1.
Baseline Rigid End Effector Operation Similarly, for the baseline end effector, we employ a $2\times 3$ grid of electrodes, shown in Fig. 2 (g), to detect the relative pose of the end effector in relation to the human limb. This end effector follows the same design and control paradigm specified in [7, 9]. Specifically, the $2$-left-middle and $5$-right-middle electrodes are essential for identifying contact (along the $z$ axis) between the baseline end effector and human skin. The four electrodes, numbered $1,4,3$ and $6$, enable the detection of displacement along the $x$ axis and the measurement of the yaw rotation $\alpha$ on the surface of the limb, ensuring that the center line of the end effector remains aligned with the central axis of a human limb. Furthermore, the two rows of six electrodes collectively facilitate the measurement of roll rotation $\gamma$ on the limb’s surface, ensuring that the bottom surface of the end effector remains parallel to human skin. Note that we do not detect translations along the $y$ axis or rotation $\beta$ around the $y$-axis, as shown in Fig. 2 (f). The end effector translates at a fixed velocity of 4 cm/s along the axis of the limb to ensure bathing task progress.
@@
Algorithm 2 defines the capacitive servoing control strategy for the baseline rigid end effector. Similar to the approach used with the SkinGrip, at each time step, we normalize all real-time capacitance values to the range $[0,1]$ using the formula $\bm{C}_{norm}^{r}=\frac{\sum\bm{C}^{r}_{t-4:t}/n-\bm{C}_{min}^{r}}{\bm{C}_{max}^{r}-\bm{C}_{min}^{r}}$, where $\bm{C}_{max}^{r}$ and $\bm{C}_{min}^{r}\in\mathbb{R}^{6}$ denote the maximum and minimum capacitance values, respectively, for each electrode. We establish a threshold $\bm{C}_{th}^{r}\in\mathbb{R}^{6}$ to ensure full contact between the baseline end effector and the human skin. The control for the baseline end effector is then defined as:
@@
where $x$ and $z$ represent the x- and z-coordinate values of the displacements in the baseline end effector frame, and $\alpha$ and $\gamma$ denote the orientation values for roll and yaw displacements in the same frame (Fig. 2 (f)). The term $\bm{K}_{p}^{r}\in\mathbb{R}^{4\times 4}$ denotes a diagonal matrix for the proportional (P) gains. Specifically, $x_{r}=(\bm{C}_{norm}^{r})_{1}-(\bm{C}_{norm}^{r})_{3}+(\bm{C}_{norm}^{r})_{4}-(\bm{C}_{norm}^{r})_{6}$ is determined by the normalized capacitance values from electrodes in both the first row (electrodes $1$ and $3$) and the second row (electrodes $4$ and $6$), which are used to measure lateral displacement along the x-axis (Fig. 2 (g)). Similarly, $z_{r}=(\bm{C}_{th}^{r})_{2}-(\bm{C}_{norm}^{r})_{2}+(\bm{C}_{th}^{r})_{5}-(\bm{C}_{norm}^{r})_{5}$ is calculated from the difference between the threshold and normalized capacitance values of electrodes $2$ and $5$, enabling the assessment of vertical distance along the $z$-axis. The orientation angles $\alpha_{r}=(\bm{C}_{norm}^{r})_{4}-(\bm{C}_{norm}^{r})_{1}+(\bm{C}_{norm}^{r})_{3}-(\bm{C}_{norm}^{r})_{6}$ and $\gamma_{r}=(\bm{C}_{norm}^{r})_{1}+(\bm{C}_{norm}^{r})_{2}+(\bm{C}_{norm}^{r})_{3}-(\bm{C}_{norm}^{r})_{4}-(\bm{C}_{norm}^{r})_{5}-(\bm{C}_{norm}^{r})_{6}$ are obtained from the configuration of six electrodes in two rows, which facilitate the detection of roll and yaw movements relative to the limb’s surface and its central axis, respectively. As with the SkinGrip, the rigid end effector follows a linear trajectory along the limb. In this process, the manipulator’s $x$, $z$, $\alpha$ and $\gamma$ positions will be updated using 3.
@@ IV Study Design
To assess the cleaning performance of both our SkinGrip and the baseline end effector, we conducted a human-robot study involving $12$ able-bodied adult participants [$2$ females, $10$ males, age: $26.9\pm 6.5$ years, weight: $73.3\pm 9.3$ kg, height: $176.6\pm 8.2$ cm]. This study was approved by Carnegie Mellon University’s Institutional Review Board under 2022.00000273.
The experimental setup is illustrated in Fig. 4 (a). Participants lie on a hospital bed with the head and foot of the bed adjustable for comfort based on individual preferences. A limb holder secures the participant’s limb parallel to the ground plane during the cleaning task, and three cameras are positioned to capture top, side, and bottom views of the participant’s limbs. Prior to cleaning, three rings of shaving cream are applied to the limb that the robot must clean off to demonstrate skin coverage and bed bathing performance. Shaving cream was selected as it is skin-safe, straightforward to apply, and has a color that contrasts with skin tones to support visual analysis in cleaning performance. Recorded images from all three cameras are used to evaluate the cleaning performance of both the soft and baseline end effectors. An additional phone camera records video throughout the experiment. For each participant and each end effector, the robot conducted 2 bathing attempts for both the arm and the leg, resulting in 8 bathing trials per participant. We randomized ordering between the two end effectors for each participant.
At the beginning of an experiment with each participant, we established specific capacitance thresholds for both end effectors. The thresholds for the SkinGrip were identified at the point where participants reported comfort and confirmed full contact between the soft finger of the end effector and their limb as the soft fingers encircled their limb. In contrast, thresholds for the rigid end effector were set based on participant reports of comfortable pressure when the end effector, held at a neutral orientation with no roll, pitch, or yaw, was placed centrally against the skin of their limb. Note that SkinGrip is designed for bed bathing rather than fully submerged or high-humidity environments such as showering. The capacitive sensors operate reliably in mild moisture conditions and are protected with a plastic film to prevent direct contact with water, the washcloth, or human skin. As a result, the system does not require full waterproofing to perform effective cleaning in bed bathing scenarios.
@@ V-C Discussion
Through this research, we have illustrated the practicality and potential benefits of our soft robotic manipulator in enhancing the quality of bathing assistance, with an emphasis on safety, comfort, and effectiveness. However, our work builds on a few assumptions that could benefit from future advances. One key assumption is the expectation that an individual will be able to position their limb static, lifted off of a bed by a limb support. Second, we evaluated the proposed SkinGrip primarily on limb bathing scenarios. Although our SkinGrip has potential to clean other body parts, such as the back and torso, the control methods required for these tasks are still necessary. While the $1$-DOF soft fingers enable effective cleaning, future work could explore higher-DOF designs to further improve adaptability on complex body contours. An additional future development may include the design of computer vision techniques to track regions of a limb that have not been fully cleaned in order to perform a targeted cleaning maneuver around those regions.
@@ VI Conclusion
In this work, we introduced an innovative soft deformable robotic manipulator—SkinGrip, designed for high adaptability and conforming to the contours of human limbs. By incorporating a capacitive servoing control scheme, the SkinGrip is able to detect contact and tightness of its soft fingers around the circumference of a human limb, while also effectively traversing the limb to perform robot-assisted bed bathing. We evaluate the SkinGrip and control strategy in a human experiment in which a mobile manipulator assists with cleaning off the arm and leg of 12 human participants. Our experiments highlight the SkinGrip and capacitive servoing control strategy’s ability to clean and adapt to various limb sizes, including differences in diameter and length. Notably, the SkinGrip achieves an overall average of $88.8\%$ cleaning effectiveness on the arm and $81.4\%$ on the leg, significantly outperforming a baseline rigid end effector design. Participant-provided feedback reinforces this finding, in which we observe a statistically significant difference in reported preferences for a soft manipulation strategy in terms of safety, comfort, and thorough skin bathing capabilities.